OpenStudy (anonymous):

New question what is the vertex of f(x)=x^2+10x+16? A)(2,6) B)(1,5) C)(5,9) D)(-5,-9)

OpenStudy (misty1212):

HHI!!!

OpenStudy (misty1212):

the first coordinate of the vertex of \(y=ax^2+bx+c\) is \(-\frac{b}{2a}\) which in your case is \[-\frac{10}{2\times 1}\]

OpenStudy (misty1212):

there is only one choice with \(-5\) as the x coordinate, so you don't have to find the second coordinate

OpenStudy (anonymous):

so the correct answer is D?

OpenStudy (misty1212):

lol yeah that is the only one with a -5 in the first slot right?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

lol just reassuring in case someone has this same question and they wnt to answer it quickly

OpenStudy (anonymous):

thank you :D!

OpenStudy (anonymous):

you welcome

OpenStudy (anonymous):

so how old are u